OpenStudy (anonymous):

(12x^2/5y) x( 9y^2/8x^6)

OpenStudy (lgbasallote):

here's a hint \[\huge \frac{12x^2}{5y} \times \frac{9y^2}{8x^6} \implies \frac{9y^2}{5y} \times \frac{12x^2}{8x^6}\] does that help?

OpenStudy (anonymous):

anything helps.thanks

OpenStudy (lgbasallote):

let me give you a hint on what my hint meant...if you look closely at \[\huge \frac{9y^2}{5y} \times \frac{12x^2}{8x^6}\] you'll see some things that will cancel out

OpenStudy (anonymous):